# Context Kubernetes: Declarative Orchestration of Enterprise Knowledge for Agentic AI Systems

### 14 Apr 2026 v2

@@ Abstract.
We introduce Context Kubernetes, an architecture for orchestrating enterprise knowledge in agentic AI systems, together with a prototype implementation and experimental evaluation. The core observation is that delivering the right knowledge, to the right agent, with the right permissions, at the right freshness, within the right cost envelope—across an entire organization—is structurally analogous to the container orchestration problem that Kubernetes solved a decade ago. We develop this analogy into six core abstractions, a YAML-based declarative manifest for knowledge-architecture-as-code, a reconciliation loop, and a three-tier agent permission model where agent authority is always a strict subset of human authority. We implement a prototype ($\sim$}7,000 lines, 92 tests) and evaluate it through eight experiments. Three value experiments provide the headline findings: (1) without governance, agents serve phantom content from deleted sources, deliver contradictory information, and leak cross-domain data in 26.5% of queries; governed routing eliminates phantom delivery and reduces noise by 14 percentage points; (2) without freshness monitoring, stale and deleted content is served silently; with reconciliation, staleness is detected in under 1 ms; (3) in five realistic attack scenarios, flat permissions block 0/5 attacks, basic RBAC blocks 4/5, and the three-tier model blocks all five—the attack RBAC misses is the one the three-tier model is specifically designed to catch. Five correctness experiments confirm: zero unauthorized context deliveries across all test queries, zero permission invariant violations, and architectural enforcement of out-of-band approval isolation that no surveyed enterprise platform provides. We identify four properties—heterogeneity, semantics, sensitivity, and learning—that make context orchestration harder than container orchestration, and argue that these properties make the solution more valuable. We conclude that Context Engineering will emerge as the defining infrastructure discipline of the AI era.
@@ 1.2. Contributions
    A prototype implementation ($\sim$}7,000 lines of Python, 92 tests) comprising Context Router, Permission Engine, CxRI connectors, reconciliation loop, audit log, and HTTP API (Section 6).
@@
    Five correctness experiments demonstrating zero unauthorized context deliveries, zero permission invariant violations, sub-millisecond freshness detection, acceptable latency overhead, and architectural enforcement of approval isolation (Section 7.1).
@@
    Three value experiments demonstrating what goes wrong without governance (phantom content, data leaks, contradictory information) and what the three-tier model catches that alternatives do not (Section 7.2).
@@ 3.4. The Agent Permission Model
This invariant is architectural: the Permission Engine computes $P_{a_{u}}$ as a configured projection of $P_{u}$. No API exists for an agent to modify its own permission profile. We note that this invariant is a design commitment, not a formally verified property; a complete implementation would require verification that no code path allows escalation, which is a standard security engineering task but is not addressed in this paper.
@@ 6. Implementation
We implement Context Kubernetes as an open-source Python prototype. [^1] Source code: https://github.com/Cohorte-ai/context-kubernetes The implementation comprises seven components totaling $\sim$}7,000 lines with 92 automated tests:
@@
The prototype uses rule-based intent classification rather than LLM-assisted parsing. This is a deliberate design choice: it establishes a performance floor that any LLM integration can only improve, and it ensures all experiments are reproducible without API dependencies. Section 9.2 discusses the implications.
Seed data for the experiments models a 10-person consulting firm with three clients, five context domains (clients, sales, delivery, HR, finance), and 12 context files containing realistic organizational knowledge.
@@ 7. Evaluation
We evaluate Context Kubernetes through eight experiments in two categories: five correctness experiments that verify the system does what it claims, and three value experiments that demonstrate why governance matters. All experiments run against the prototype with the seed dataset.
@@ 7.1. Correctness Experiments
C1: Routing quality. 47 queries across 5 domains with ground-truth relevance labels. The rule-based router achieves 70.2% domain classification accuracy and 0.665 MRR (Mean Reciprocal Rank). Average routing latency: 202 ms. This establishes a floor: LLM-assisted parsing would improve domain accuracy but the permission and freshness guarantees hold regardless of classification quality.
@@
7.2. Value Experiments
The correctness experiments show the system works. The value experiments show why it matters—what goes wrong without governance and what the three-tier model catches that alternatives do not.
V1: Governed vs. ungoverned context quality. The same 47 queries are processed through two pipelines: (a) ungoverned naive RAG (embed query, cosine similarity, top-$k$, no permissions, no freshness) and (b) Context Kubernetes (intent routing, permission filtering, freshness filtering, multi-signal ranking). Results:
Table 6. Governed vs. ungoverned context delivery (47 queries)
Metric Ungoverned Governed
Content leak rate (cross-domain) 26.5% 21.6%
Noise ratio (irrelevant units delivered) 81.6% 67.6%
Total content leaks 88 76
Token waste 80.0% 79.9%
The governed pipeline reduces noise by 14 percentage points and eliminates the most egregious leaks (HR data to sales agents, finance data to unauthorized roles). The modest improvement in token waste reflects the limitation of rule-based routing; LLM-assisted intent parsing would narrow the token budget to truly relevant content. The content leak reduction from 88 to 76 is driven entirely by the permission filter—the router retrieves similar candidates, but the Permission Engine strips unauthorized units before delivery.
@@ 8. Related Work
Role-based access control. Sandhu et al. (Sandhu et al., 1996) formalized RBAC for human identities. Context Kubernetes extends RBAC to agent identities with the strict-subset invariant (Design Invariant 3.7) and three-tier approval. This extension is, to our knowledge, novel; we are not aware of prior RBAC formalizations that model autonomous agents as permission-subset principals of their human operators.
@@ 9.1. Limitations
Rule-based routing only. The prototype uses keyword-based intent classification, not LLM-assisted parsing. This establishes a quality floor (70.2% domain accuracy, 0.665 MRR) but understates the architecture’s potential. The governed pipeline’s modest precision improvement over ungoverned RAG (Table 6) is primarily a limitation of rule-based classification, not of governance itself—the permission and freshness layers deliver their guarantees regardless of routing quality.
@@
Design invariants tested, not formally proven. The 92 tests verify invariant enforcement empirically (including adversarial attempts). TLA+ and Alloy specifications have been written but not yet model-checked. We use “design invariant” and “design goal” to distinguish from formally verified properties.
@@ 9.2. The LLM-in-the-Loop Tension
None of these fully resolves the tension. We consider this an open problem: how to build governance-grade infrastructure that incorporates probabilistic AI components without undermining its own guarantees. This is likely a general challenge for AI-native systems, not specific to Context Kubernetes.
@@ 9.6. Future Work
Three directions would most strengthen the contribution. First, LLM-assisted intent routing: the prototype’s rule-based classifier is a floor; integrating a small model (e.g., gpt-4o-mini) for semantic intent parsing would improve domain accuracy and relevance precision, making the governed pipeline’s quality advantage over naive RAG more dramatic. Second, production deployment: a 4-week pilot at a real organization would replace seed-data experiments with real-world metrics and surface design problems that lab conditions cannot. Third, formal verification: the TLA+ and Alloy specifications are written; model-checking them would elevate the safety and liveness claims from “tested empirically” to “verified formally.”
@@ 10. Conclusion
The experiments demonstrate three things. First, governance matters: without it, agents serve phantom content, deliver contradictory information, and leak cross-domain data—problems that are invisible until they cause damage. Second, the three-tier permission model catches what RBAC cannot: in five realistic attack scenarios, RBAC blocks four but misses the one where an agent sends confidential pricing via email—the exact scenario that motivates separating agent authority from user authority. Third, the architecture’s safety and freshness guarantees hold: zero unauthorized context deliveries, sub-millisecond staleness detection, and architectural enforcement of approval isolation that no surveyed enterprise platform provides.
@@
The most consequential claim in this paper is not architectural but organizational: Context Engineering will emerge as the defining infrastructure discipline of the AI era, just as DevOps emerged for the container era. The organizations that invest in context infrastructure—declarative knowledge architectures, governed routing, organizational intelligence operators—will compound their advantage with every deployment.

### 16 Apr 2026 v3

@@ Abstract.
We introduce Context Kubernetes, an architecture for orchestrating enterprise knowledge in agentic AI systems, together with a prototype implementation and experimental evaluation. The core observation is that delivering the right knowledge, to the right agent, with the right permissions, at the right freshness, within the right cost envelope—across an entire organization—is structurally analogous to the container orchestration problem that Kubernetes solved a decade ago. We develop this analogy into six core abstractions, a YAML-based declarative manifest for knowledge-architecture-as-code, a reconciliation loop, and a three-tier agent permission model where agent authority is always a strict subset of human authority. We implement a prototype (92 automated tests) and evaluate it through eight experiments. Three value experiments provide the headline findings: (1) across 200 benchmark queries on synthetic seed data, we compare four governance baselines from ungoverned RAG to ACL-filtered retrieval to RBAC-aware routing to the full architecture; each layer contributes a different capability, and only the three-tier model blocks all five tested attack scenarios; (2) without freshness monitoring, stale and deleted content is served silently; with reconciliation, staleness is detected in under 1 ms; (3) in five realistic attack scenarios, flat permissions block 0/5 attacks, basic RBAC blocks 4/5, and the three-tier model blocks all five—the attack RBAC misses is the one the three-tier model is specifically designed to catch. Five correctness experiments confirm: zero unauthorized context deliveries, zero permission invariant violations, and architectural enforcement of out-of-band approval isolation that no surveyed enterprise platform provides. TLA+ model-checking verifies the safety properties across 4.6 million reachable states with zero violations. We identify four properties—heterogeneity, semantics, sensitivity, and learning—that make context orchestration harder than container orchestration, and argue that these properties make the solution more valuable. We conclude that Context Engineering will emerge as the defining infrastructure discipline of the AI era.
@@ 1.2. Contributions
    A prototype implementation (92 tests) comprising Context Router, Permission Engine, CxRI connectors, reconciliation loop, audit log, and HTTP API (Section 6).
@@
    Five correctness experiments demonstrating zero unauthorized context deliveries, zero permission invariant violations, sub-millisecond freshness detection, acceptable latency overhead, and architectural enforcement of approval isolation—with safety properties formally verified by TLC model-checking across 4.6 million reachable states (Section 7.1).
@@
    Three governance impact experiments on synthetic data demonstrating what goes wrong without governance (phantom content, data leaks, contradictory information) and what the three-tier model catches that alternatives do not (Section 7.2).
@@ 3.4. The Agent Permission Model
This invariant is enforced at registration time: the Permission Engine rejects any agent profile that is not a strict subset, that includes operations outside the user’s role, or that specifies a less restrictive tier than the user’s. The 92 tests include adversarial attempts to violate the invariant (superset registration, equal-set registration, less-restrictive tier), all of which are correctly rejected. An Alloy specification encoding the invariant as five formal assertions has been written (formal/PermissionModel.als) but has not yet been model-checked. The specification is complete and committed to the public repository; we retain “design invariant” rather than “verified property” pending execution of the Alloy Analyzer, which is a GUI-based tool requiring manual operation. If the Alloy check found a counterexample, it would indicate a configuration in which the strict-subset property can be violated despite the registration-time enforcement—a finding that would require adding runtime permission checks in addition to the current registration-time guards.
@@ 6. Implementation
We implement Context Kubernetes as an open-source Python prototype. [^1] Source code: https://github.com/Cohorte-ai/context-kubernetes The implementation comprises seven components with 92 automated tests:
@@
The prototype supports both rule-based and LLM-assisted (gpt-4o-mini) intent classification. Rule-based classification serves as the deterministic floor (63.0% domain accuracy), ensuring all experiments are reproducible without API dependencies. LLM-assisted classification improves domain accuracy to 75.0% and is used in the value experiments to demonstrate the quality ceiling. Section 9.2 discusses the tension between probabilistic routing and governance guarantees.
Seed data for the experiments models a 10-person consulting firm with three clients, five context domains (clients, sales, delivery, HR, finance), and 12 context files containing realistic organizational knowledge. The benchmark comprises 200 queries with ground-truth relevance labels (50 sales, 40 delivery, 30 HR, 30 finance, 50 cross-domain).
@@ 7. Evaluation
We evaluate Context Kubernetes through eight experiments in two categories: five correctness experiments that verify the system does what it claims, and three governance impact experiments that demonstrate what goes wrong without governance. All experiments run against the prototype with the seed dataset.
@@ 7.1. Correctness Experiments
C1: Routing quality. 200 queries across 5 domains with ground-truth relevance labels, tested with both rule-based and LLM-assisted (gpt-4o-mini) intent classification. Rule-based: 63.0% domain accuracy, 162 ms average latency. LLM-assisted: 75.0% domain accuracy (+12 pp), 1,589 ms average latency (dominated by API call). The permission and freshness guarantees hold regardless of classification mode or accuracy—a 25% misrouting rate means 25% of queries receive irrelevant results, but zero queries receive unauthorized results. The governance layer operates below and independently of the router: it does not trust the domain classification. Governance guarantees are therefore accuracy-independent. However, practical utility is not: at 75% routing accuracy, one in four queries returns irrelevant (though never unauthorized) results. An enterprise deployment would likely require $\geq$90% routing accuracy for the system to deliver value, even though its safety guarantees hold at any accuracy including zero.
@@
7.2. Governance Impact Experiments
The correctness experiments show the system works as designed. The governance impact experiments show what goes wrong without governance and what the three-tier model catches that alternatives do not. All experiments run on synthetic seed data designed by the author; the results demonstrate that the governance mechanisms function correctly, not that they have been validated on real organizational data. Validation on naturally occurring enterprise queries is the critical next step (Section 9.6).
V1: Governed vs. ungoverned context quality. 200 queries are processed through four pipelines of increasing governance strength, including two intermediate baselines requested by reviewers to ensure the comparison is not stacked in favor of the proposed system:
Table 6. Four governance baselines compared (200 queries, synthetic data)
Baseline Leaks Leak% Noise% Attacks
B0: Ungoverned RAG 375 26.5% 80.9% 0/5
B1: ACL-filtered RAG 0 0.0% 82.8% 4/5
B2: RBAC-aware RAG 295 20.9% 64.0% 4/5
B3: Context Kubernetes 307 20.9% 64.0% 5/5
B0: cosine top-$k$, no permissions, no freshness. B1: B0 + post-retrieval role-based document filtering. B2: intent routing + user-role permissions applied to agent (no agent-specific subset, no tiered approval). B3: full Context Kubernetes (intent routing + agent permission subset + three-tier approval + freshness). LLM-assisted routing (not shown) reduces B3 noise from 64.0% to 35.1% and leaks from 307 to 161.
The comparison reveals that each governance layer contributes a different capability. ACL filtering (B1) eliminates cross-domain leaks entirely—a strong baseline that any enterprise should implement as a minimum. Intent routing (B2, B3) reduces noise by 19 percentage points (82.8% $\to$ 64.0%) by directing queries to relevant domains rather than searching all content. The three-tier model (B3 only) is the sole approach that blocks all five attack scenarios—the one attack B1 and B2 miss is the agent sending confidential pricing via email, because RBAC allows the role to send email and cannot distinguish email-with-pricing from email-without-pricing. The three-tier model can, because agent-level operation classification is decoupled from role-level authorization.
Context Kubernetes does not claim to outperform every alternative on every metric. ACL-filtered retrieval achieves zero leaks, which B3 does not (because B3 enables cross-domain brokering, which increases completeness at the cost of a larger authorization surface). The unique contribution of the full stack is the combination of noise reduction, freshness governance, and tiered approval—capabilities that no single simpler baseline provides together.
@@ 8. Related Work
Role-based access control and capability-based security. Sandhu et al. (Sandhu et al., 1996) formalized RBAC for human identities. The strict-subset containment principle has precedent in capability-based security: the Principle of Least Authority (POLA) in operating systems research, seL4’s formally verified domain separation, and SOAAP’s compartmentalization model all enforce the property that a delegated process cannot exceed its delegator’s authority. Context Kubernetes applies this principle to a setting these systems did not address: autonomous AI agents that can reason, plan, and attempt multi-step circumvention, acting as delegated principals of human users. The three-tier approval model (particularly the out-of-band Tier 3 isolation) is, to our knowledge, novel in this agent-delegation context.
@@ 9.1. Limitations
Routing quality depends on classification mode. The prototype supports both rule-based (63.0% domain accuracy) and LLM-assisted (75.0%) intent classification. LLM-assisted governance reduces leaks by 57% compared to ungoverned delivery, but adds $\sim$}1.4 seconds of latency per query due to the API call. A production system would need a local model or a faster API to achieve both quality and latency targets simultaneously.
@@
Safety formally verified; liveness empirically tested. The TLA+ specification of the reconciliation loop has been model-checked with TLC across 4,615,030 distinct reachable states (33.6 million states generated, search depth 30). Safety properties S1 (no expired content served) and S2 (fail-closed when Permission Engine is down) hold in every reachable state—zero violations. Liveness properties (stale detection within bounded time) are verified empirically by the prototype’s 92 tests, including 11 reconciliation tests with sub-millisecond detection latency. The Alloy permission model specification is structurally complete; its five assertions (Invariants 3.7a/b/c, 3.8, and no-escalation) follow logically from the declared axioms and are expected to pass automated checking. We retain “design invariant” for the permission model until Alloy verification is executed.
@@ 9.2. The LLM-in-the-Loop Tension
None of these fully resolves the tension. However, a critical architectural property limits the blast radius of misrouting: routing misclassification cannot produce permission leaks (given the deployment preconditions in Section 9.4, particularly that CxRI connectors are not compromised). The Permission Engine operates below the router — it filters every returned Context Unit by the requesting agent’s authorized_roles, regardless of which domain the router selected. A query misrouted to the wrong domain returns irrelevant results (a quality failure) but cannot return unauthorized results (a security failure). This separation is architectural: the Permission Engine does not trust or depend on the router’s domain classification. The safety guarantees verified by TLC (no expired content served, fail-closed on Permission Engine down) hold regardless of routing accuracy. Misrouting degrades relevance; it does not degrade governance.
We consider the broader tension an open problem: how to build governance-grade infrastructure that incorporates probabilistic AI components without undermining its own guarantees. This is likely a general challenge for AI-native systems, not specific to Context Kubernetes.
@@ 9.6. Future Work
Two directions would most strengthen the contribution. First, production deployment: a 4-week pilot at a real organization would replace seed-data experiments with real-world metrics and surface design problems that lab conditions cannot. Second, completing formal verification: TLA+ safety properties are verified; the remaining steps are Alloy model-checking for the permission invariants and formal liveness verification (which requires more sophisticated temporal reasoning than bounded model-checking supports).
@@ 10. Conclusion
The experiments demonstrate three things. First, governance is layered: across 200 benchmark queries and four baselines of increasing strength, ACL filtering eliminates cross-domain leaks, intent routing reduces noise by 19 percentage points, and only the three-tier model blocks all five tested attack scenarios. No single simpler alternative provides all three capabilities. Second, the three-tier permission model catches what RBAC cannot: in five realistic attack scenarios, RBAC blocks four but misses the one where an agent sends confidential pricing via email—the exact scenario that motivates separating agent authority from user authority. Third, the architecture’s safety guarantees hold—not just empirically but formally: TLC model-checking verifies across 4.6 million reachable states that no expired content is ever served and the system always fails closed when the Permission Engine is unavailable. Zero unauthorized context deliveries, sub-millisecond staleness detection, and architectural enforcement of approval isolation that no surveyed enterprise platform provides.
@@
The most consequential claim in this paper is not architectural but organizational: we argue that Context Engineering will emerge as the defining infrastructure discipline of the AI era, just as DevOps emerged for the container era. The organizations that invest in context infrastructure—declarative knowledge architectures, governed routing, organizational intelligence operators—will compound their advantage with every deployment.